OpenStudy (anonymous):

i WILL GIVE YOU A MEDAL AND BE YOU FAN!! What is 7/10 divided by 4/5? A. 14/25 B. 1 1/7 C. 7/8 D.1 11/14

OpenStudy (solomonzelman):

dividing by 4/5 is same as multiplying times 5/4.

OpenStudy (hoblos):

\[\frac{ a }{ b } \div \frac{ c }{ d } = \frac{ a }{ b }\times \frac{ d }{ c }\]

OpenStudy (anonymous):

c 7/8 = 7/10 x 5/4

OpenStudy (anonymous):

This dosen't make any sense

OpenStudy (hoblos):

\[\frac{ 7 }{ 10 }\div \frac{ 4 }{ 5 } = \frac{ 7 }{ 10 }\times \frac{ 5 }{ 4 } = \frac{ 35 }{ 40 } = \frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

Ok,now I understand it!! Thanks

OpenStudy (hoblos):

any time :)